You are conducting a study to see if the accuracy rate for fingerprint identification is significantly more than 0.88. You use a significance level of α=0.01α=0.01.

      H0:p=0.88H0:p=0.88
      H1:p>0.88H1:p>0.88

You obtain a sample of size n=587 in which there are 528 successes.

What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic =

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value =

The p-value is...

a) less than (or equal to) αα

b) greater than αα



This test statistic leads to a decision to...

a) reject the null

B) accept the null

C) fail to reject the null



As such, the final conclusion is that...

a) There is sufficient evidence to warrant rejection of the claim that the accuracy rate for fingerprint identification is more than 0.88.

b) There is not sufficient evidence to warrant rejection of the claim that the accuracy rate for fingerprint identification is more than 0.88.

c) The sample data support the claim that the accuracy rate for fingerprint identification is more than 0.88.

d) There is not sufficient sample evidence to support the claim that the accuracy rate for fingerprint identification is more than 0.88.
Answer #1

Test hypothesis H, p = 0.88 H p088 (claim) Given sample size n=587,X-528 Sample proportion52-0.899489 Test statistic z = (1-p)/n n 587 p-p 0.899489-0.88 0.88(1-0.88)/587 = 1.453 P-value P(1.453) = 1-P(zsl .453) 1-0.9269 -0.0731The p-value is Since p-value>0.01 (a) Option b is correct The test statistic leads to a decision to Fail to reject the null hypothesis Option c is corTect Conclusion There is not sufficient evidence to support the claim that the accuracy rate for fingerprint identification is more than 0.88 Option d is correct